An unknown oxalate, XC2O4, is found to be 41.72% O by mass. Identify the element, X.


1
Expert's answer
2021-02-28T06:19:28-0500

Solution:

Suppose the molar mass of X is x, then:

M(XC2O4) = Ar(X) + 2×Ar(C) + 4×Ar(O) = x + 2×12.0107 + 4×15.999 = x + 88.0174 (g/mol)

Hence,

w(O) = [4×Ar(O) / M(XC2O4)] × 100%

0.4172 = 4×15.999 / (x + 88.0174)

x + 88.0174 = 153.394

x = 65.376 (g/mol)

The molar mass equal to 65.376 g/mol corresponds to the element - zinc (Zn).

Thus: XC2O4 = ZnC2O4 - zinc oxalate


Answer: Element X is zinc (Zn).